OpenStudy (anonymous):

police lieutenants, examining the scene of an accident involving two cars, measure the skid marks of one of the cars, which nearly came to a stop before colliding, to be 80m long. the coefficient of kinetic friction between rubber and the pavement is about 0.80. estimate the initial speed of that car assuming a level road. police lieutenants, examining the scene of an accident involving two cars, measure the skid marks of one of the cars, which nearly came to a stop before colliding, to be 80m long. the coefficient of kinetic friction between rubber and the pavement is about 0.80. estimate the

OpenStudy (fretje):

I calculate the time interval between the beginning of slipping, and the end of slipping (collision): s=a.t²/2 and a = F/masscar and F = Phi = Muk.N (Muk is friction coeff 0,80) and N = masscar.g Solving for t gives t ²= 20.38736 thus t = 4.515 seconds (4 seconds and a half approx)

OpenStudy (fretje):

then v0 = a.t = 2.ds/t with ds distance travelled (80m) ds = a.t²/2 => a = 2*80m/20,38736 = 7.848 m/s² V0 = a.t = 35,43m/s